OpenStudy (anonymous):

Help with this equation, show your work please!! 1/a=6/18

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

1/a=1/3 so a=3

OpenStudy (anonymous):

a*1/a=6/18 *a (multiply a to both sides) 1 *18= 6a/18 * 18 (multiply 18 to both sides) 18 /6 = 6a /6 (divide 6 to both sides ) 3 = a